A 20kg box is at rest on a friction less surface.then,a man pushed the box to a velocity of 3 m/s for 2 s


1
Expert's answer
2021-06-07T09:52:55-0400

The given surface is friction less and initial velocity of the block is 0(rest).

Acceleration is change in velocity per unit time.


"acceleration(a)= \frac{\triangle V}{\triangle t} \newline
a=\frac{3}{2}\newline
acceleration = 1.5 \ m/s^2"

Thus the force applied on the box by the man = m*a

where m- mass of box and a- acceleration.


"Force(F)=Mass(m)\times Acceleration(a) \newline
F=20 \times1.5 \newline
F=30 \ N"